Find the molecular formula of nicotine given the empirical formula C5H7N wuth the molecular mass of 194.
1
Expert's answer
2021-10-08T02:07:46-0400

"C_5H_7N \to 5\times 12+ 7\times 1 +14=81"

Molecular mass =194

"n=194\/81=2.425=2"

molecular formula "=n \times" empirical formula"=2\times C_5H_7N=C_{10}H_{14}N_2"